OpenStudy (anonymous):

Solve using the substitution method. 1/3x-1/5y= 7 1/6x+2/3= -4

OpenStudy (amistre64):

i dont see why we need to be constrained to the substitution method ....

OpenStudy (anonymous):

Solve the second equation for X and then plug that value in for the X in the first equation.

OpenStudy (anonymous):

I'm assuming it wants the Y-value?

OpenStudy (anonymous):

To use the substitution method, solve the second equation for x, then substitute the result in for x in the first equation and solve it for y.

OpenStudy (anonymous):

just help plz

OpenStudy (amistre64):

1/3x-1/5y= 7 ; *-1/2 1/6x+2/3y= -4 -1/6x+1/10y= -7/2 1/6x+2/3y= -4 ------------------ y(1/10+2/3) = (-4 -7/2) y = (-15/2)/(1/10+2/3) :)

OpenStudy (anonymous):

Amistre, There is no Y in the second equation.

OpenStudy (anonymous):

can u make it clear

OpenStudy (anonymous):

it's pose to be a y

OpenStudy (amistre64):

y = (-15/2)/(23/30) y = -15(30)/2(23) = -15(15)/(23) = -225/23 if i kept the fraction in line

OpenStudy (anonymous):

is the answer right

OpenStudy (amistre64):

as far as i can tell; that would be the answer for y ... who conjured up this twisted problem?

OpenStudy (anonymous):

What amistre did, was; Mutiplied each side of the first equation by -1/2. You then combine the first and second equations using addition, the x terms cancel and you're left with y = (-15/2)/(1/10+2/3) = -225/23

OpenStudy (amistre64):

colving for x is just as much a pain even when you know the value of y if you have only one option with a y=-225/23 id take it :)